OpenStudy (anonymous):

a perpendicular line is drawn from focus F of the parabola x^2 =4ay to the point X on the tangent at P(2ap, ap^2). Find the locus of X

ganeshie8 (ganeshie8):

lets try

ganeshie8 (ganeshie8):

lets call the point on tangent : X (x, y)

ganeshie8 (ganeshie8):

x^2=4ay, can we write the tangent equation at P(2ap, ap^2) ?

OpenStudy (anonymous):

yes just a sec

ganeshie8 (ganeshie8):

okayy

OpenStudy (anonymous):

y-ap^2 =p(x-2ap)

ganeshie8 (ganeshie8):

good :) calculate the distance from Focus to the tangent

ganeshie8 (ganeshie8):

use the same formula we used last time : perpendicular distance form Focus to the tangent line : (Ax1+BY1+C)/sqrt(A^2+B^2)

ganeshie8 (ganeshie8):